OpenStudy (anonymous):

Given the function f(x) = 4(2)x, Section A is from x = 1 to x = 2 and Section B is from x = 3 to x = 4. Part A: Find the average rate of change of each section. Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other.

OpenStudy (anonymous):

rate of change section A f(2) - f(1) divided by (2 - 1 ) = rate of change section B = f(4) - f(3) divided by (4 - 3) = part B section B: section A = 16/8 = 2 times

OpenStudy (anonymous):

thanks!